# Multi-Derivative Runge-Kutta Flux Reconstruction for hyperbolic conservation laws

### v1

@@ 5.2 Stage 2
$\mathsf{F}^{2}_{e}=a\mathsf{U}_{e}^{2},\qquad\mathsf{U}_{e}^{2}=\mathsf{u}_{e}^{n}-\frac{1}{6}\sigma\mathsf{D}\mathsf{u}_{e}^{n}-\frac{1}{3}\sigma\mathsf{D}\mathsf{u}_{e}^{\ast}=\mathsf{T}^{2}\mathsf{u}_{e}^{n}+\mathsf{T}^{2,\ast}\mathsf{u}_{e}^{\ast}$
@@
$\mathsf{T}^{2}=\mathsf{I}-\frac{1}{6}\sigma\mathsf{D},\qquad\mathsf{T}^{2,\ast}=-\frac{1}{3}\sigma\mathsf{D}$
@@
$\mathsf{F}_{e+\frac{1}{2}}$ $=\frac{1}{2}\left[\mathsf{V}_{R}^{T}\mathsf{F}_{e}^{2}+\mathsf{V}_{L}^{T}\mathsf{F}_{e+1}^{2}\right]-\frac{1}{2}a\left(\mathsf{V}_{L}^{T}\mathsf{U}^{2}_{e+1}-\mathsf{V}_{R}^{T}\mathsf{U}^{2}_{e}\right)$
$=\frac{1}{2}a\left[\mathsf{V}_{R}^{T}\mathsf{U}^{2}_{e}+\mathsf{V}_{L}^{T}\mathsf{U}^{2}_{e+1}\right]-\frac{1}{2}a\left(\mathsf{V}_{L}^{T}\mathsf{U}^{2}_{e+1}-\mathsf{V}_{R}^{T}\mathsf{U}^{2}_{e}\right)$
$=a\mathsf{V}_{R}^{T}\mathsf{U}_{e}^{2}$
$\mathsf{F}_{e-\frac{1}{2}}$ $=a\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{2}$
@@
$F_{h}^{\delta}(x_{e+\frac{1}{2}}^{-})$ $=\mathsf{V}_{R}^{T}\mathsf{F}_{e}^{2}=a\mathsf{V}_{R}^{T}\mathsf{U}_{e}^{2}$
$F_{h}^{\delta}(x_{e-\frac{1}{2}}^{+})$ $=\text{{{V}}}_{L}^{T}\mathsf{F}_{e}^{2}=a\mathsf{V}_{L}^{T}\mathsf{U}_{e}^{2}$
@@
$\begin{split}F_{e+\frac{1}{2}}-F_{h}^{\delta}(x_{e+\frac{1}{2}}^{-})&=0\\ F_{e-\frac{1}{2}}-F_{h}^{\delta}(x_{e-\frac{1}{2}}^{+})&=a\left(\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{2}-\mathsf{V}_{L}^{T}\mathsf{U}_{e}^{2}\right)\end{split}$ (19)
@@
$\begin{split}\partial_{\xi}\mathsf{F}_{h}&=a\mathsf{D}\mathsf{U}^{2}_{e}+a\mathsf{b}_{L}\left(\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{2}-\mathsf{V}_{L}^{T}\mathsf{U}_{e}^{2}\right)\\ &=a\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{2}+a\left(\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\right)\mathsf{U}^{2}_{e}\end{split}$ (20)
We now expand $\mathsf{U}_{e}^{2}$ in terms of $\mathsf{u}_{e}^{n}$ as follows
$\mathsf{U}_{e}^{2}=\mathsf{T}^{2}\mathsf{u}_{e}^{n}+\mathsf{T}^{2,\ast}\mathsf{u}_{e}^{\ast}$
where
$\mathsf{T}^{2}=\mathsf{I}-\frac{1}{6}\sigma\mathsf{D},\qquad\mathsf{T}^{2,\ast}=-\frac{1}{3}\sigma\mathsf{D}$
Thus, by
$\mathsf{u}^{n+1}=\mathsf{u}^{n}-\frac{\Delta t/2}{\Delta x_{e}}\partial_{\xi}\mathsf{F}_{h}$
@@
$\begin{split}\mathsf{A}_{-2}&=-\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{T}^{2,\ast}\mathsf{A}^{\ast}_{-1}\\ \mathsf{A}_{-1}&=\mathsf{b}_{L}\mathsf{V}_{R}^{T}\left(\mathsf{T}^{2}+\mathsf{T}^{2,\ast}\left(1-\sigma\mathsf{A}_{0}^{\ast}\right)\right)-\sigma\left(\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\right)\mathsf{T}^{2,\ast}\mathsf{A}^{\ast}_{-1}\\ \mathsf{A}_{0}&=-\left(\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\right)\left(\mathsf{T}^{2}+\mathsf{T}^{2,\ast}\left(1-\sigma\mathsf{A}_{0}^{\ast}\right)\right)\\ \mathsf{A}_{+1}=\mathsf{A}_{+2}&=0\end{split}$ (21)
where $\mathsf{A}_{i}^{\ast}$ are defined in (18). The upwind character of D2 flux is reason why we have $\mathsf{A}_{+1}=\mathsf{A}_{+2}=0$.
@@
$\begin{array}[]{c}\widehat{\mathsf{u}}_{k}^{n+1}=H(\sigma,k)\widehat{\mathsf{u}}_{k}^{n}\\ \text{{{H}}}=-\sigma^{2}\mathsf{A}_{-2}\exp(-2i\kappa)-\sigma\mathsf{A}_{-1}\exp(-i\kappa)+\mathsf{I}-\sigma\mathsf{A}_{0}-\sigma\mathsf{A}_{+1}\exp(i\kappa)-\sigma^{2}\mathsf{A}_{+2}\exp(2i\kappa)\end{array}$
@@
The values of $\sigma$ are also sampled in some interval $[\sigma_{\min},\sigma_{\max}]$ and the largest value of $\sigma_{l}\in[\sigma_{\min},\sigma_{\max}]$ for which $\lambda(\sigma_{l})\leq 1$ is determined in a Python code. We start with a large interval $[\sigma_{\min},\sigma_{\max}]$ and then progressively reduce the size of this interval so that the CFL number is determined to about three decimal places. The results of this numerical investigation of stability are shown in Table 1 for two correction functions with polynomial degree $N=3$. The comparison is made with CFL numbers of MDRK-D1 (12) which are experimentally obtained from the linear advection test case (Section 7.1.1), i.e., using time step size that is slightly larger than these numbers causes the solution to blow up.
@@
Radau $\approx$0.09 0.107 1.19
$g_{2}$approx0.16 0.224 1.4
@@
We see that dissipation model D2 has a higher CFL number compared to dissipation model D1. The CFL numbers for the $g_{2}$ correction function are also significantly higher than those for the Radau correction function. The optimality of these CFL numbers has been verified by experiment on the linear advection test case (Section 7.1.1), i.e., the solution eventually blows up if the time step is slightly higher than what is allowed by the CFL condition.
@@ 6 Blending scheme
The MDRK scheme (7) gives a high (fourth) order method for smooth problems, but most practical problems involving hyperbolic conservation laws consist of non-smooth solutions like shocks. In such situations, using a higher order method is bound to produce Gibbs oscillations, as is stated in Godunov’s order barrier theorem [14]. The cure is to non-linearly add dissipation in regions where the solution is non-smooth, with methods like artificial viscosity, limiters and switching to a robust lower order scheme; the resultant scheme will be non-linear even for linear equations. In this work, we develop a subcell based blending scheme for MDRK similar to the one in [4] for the high order single stage Lax-Wendroff methods. The limiter is applied to each MDRK stage.
@@
$\mathsf{u}^{H,\ast}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t/2}{\Delta x_{e}}\mathsf{R}^{\ast,H}_{e},\qquad\mathsf{u}^{H,n+1}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t}{\Delta x_{e}}\mathsf{R}^{H}_{e}$ (22)
@@
$\mathsf{u}^{L,\ast}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t/2}{\Delta x_{e}}\mathsf{R}^{\ast,L}_{e},\qquad\mathsf{u}^{L,n+1}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t}{\Delta x_{e}}\mathsf{R}^{L}_{e}$
@@
$\begin{split}\mathsf{u}^{\ast}_{e}&=(1-\alpha_{e})\mathsf{u}^{H,\ast}_{e}+\alpha_{e}\mathsf{u}^{L,\ast}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t/2}{\Delta x_{e}}[(1-\alpha_{e})\mathsf{R}^{\ast,H}_{e}+\alpha_{e}\mathsf{R}^{\ast,L}_{e}]\\ \mathsf{u}^{n+1}_{e}&=(1-\alpha_{e})\mathsf{u}^{H,n+1}_{e}+\alpha_{e}\mathsf{u}^{L,n+1}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t}{\Delta x_{e}}[(1-\alpha_{e})\mathsf{R}^{n+1,H}_{e}+\alpha_{e}\mathsf{R}^{n+1,L}_{e}]\end{split}$ (23)
where $\alpha_{e}\in[0,1]$ must be chosen based on a local smoothness indicator. If $\alpha_{e}=0$ then we obtain the high order MDRK scheme, while if $\alpha_{e}=1$ then the scheme becomes the low order scheme that is less oscillatory. In subsequent sections, we explain the details of the lower order scheme and the design of smoothness indicators. The lower order scheme will either be a first order finite volume scheme or a high resolution scheme based on MUSCL-Hancock idea. In either case, the common structure of the low order scheme can be explained as follows. In the first stage, the lower order residual $\mathsf{R}^{\ast,L}_{e}$ performs evolution from time $t^{n}$ to $t^{n+\frac{1}{2}}$ while, in the second stage, $\mathsf{R}^{L}_{e}$ performs evolution from $t^{n}$ to $t^{n+1}$. It may seem more natural to evolve from $t^{n+\frac{1}{2}}$ to $t^{n+1}$ in the next stage, but that will violate the conservation property because of the expression of second stage of MDRK (522).
@@
$\begin{split}\bm{u}_{e,0}^{L,n+1}&=\bm{u}_{e,0}^{n}-\frac{\Delta t}{w_{0}\Delta x_{e}}[\bm{f}_{\frac{1}{2}}^{e}-\bm{F}_{e-\frac{1}{2}}]\\ \bm{u}_{e,p}^{L,n+1}&=\bm{u}_{e,p}^{n}-\frac{\Delta t}{w_{p}\Delta x_{e}}[\bm{f}_{p+\frac{1}{2}}^{e}-\bm{f}_{p-\frac{1}{2}}^{e}],\qquad 1\leq p\leq N-1\\ \bm{u}_{e,N}^{L,n+1}&=\bm{u}_{e,N}^{n}-\frac{\Delta t}{w_{N}\Delta x_{e}}[\bm{F}_{e+\frac{1}{2}}-\bm{f}_{N-\frac{1}{2}}^{e}]\end{split}$ (25)
The inter-element fluxes $\bm{F}_{e+\frac{1}{2}}$ used in the low order scheme are same as those used in the high order MDRK scheme in equation (6). Usually, Rusanov’s flux [25] will be used for the inter-element fluxes and in the lower order scheme. The element mean value obtained by the low order scheme satisfies
$\overline{\bm{u}}_{e}^{L,n+1}=\sum_{p=0}^{N}\bm{u}_{e,p}^{L,n+1}w_{p}=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}_{e+\frac{1}{2}}-\bm{F}_{e-\frac{1}{2}})$ (26)
@@
$=(1-\alpha_{e})\left[\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}_{e+\frac{1}{2}}-\bm{F}_{e-\frac{1}{2}})\right]+\alpha_{e}\left[\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}_{e+\frac{1}{2}}-\bm{F}_{e-\frac{1}{2}})\right]$
$=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}_{e+\frac{1}{2}}-\bm{F}_{e-\frac{1}{2}})$
@@
$\begin{split}\overline{\bm{u}}^{\ast}&=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}^{\ast}_{e+\frac{1}{2}}-\bm{F}^{\ast}_{e-\frac{1}{2}})\\ \overline{\bm{u}}_{e}^{n+1}&=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}_{e+\frac{1}{2}}-\bm{F}_{e-\frac{1}{2}})\end{split}$ (27)
@@
The inter-element fluxes $\bm{F}^{\ast}_{e+\frac{1}{2}},\bm{F}_{e+\frac{1}{2}}$ are used both in the low and high order schemes. To achieve high order accuracy in smooth regions, this flux needs to be high order accurate, however it may produce spurious oscillations near discontinuities when used in the low order scheme. A natural choice to balance accuracy and oscillations is to take
$\begin{split}\bm{F}_{e+\frac{1}{2}}&=(1-\alpha_{e+\frac{1}{2}})\bm{F}^{\operatorname{HO}\ast}_{e+\frac{1}{2}}+\alpha_{e+\frac{1}{2}}\bm{f}_{e+\frac{1}{2}}\\ \bm{F}_{e+\frac{1}{2}}&=(1-\alpha_{e+\frac{1}{2}})\bm{F}_{e+\frac{1}{2}}^{\operatorname{HO}}+\alpha_{e+\frac{1}{2}}\bm{f}_{e+\frac{1}{2}}\end{split}$ (28)
where $\bm{F}^{\operatorname{HO}\ast}_{e+\frac{1}{2}},\bm{F}_{e+\frac{1}{2}}^{\operatorname{HO}}$ are the high order inter-element time-averaged numerical fluxes used in the MDRK scheme (13) and $\bm{f}_{e+\frac{1}{2}}$ is a lower order flux at the face $x_{e+\frac{1}{2}}$ shared between FR elements and subcells (2930). The blending coefficient $\alpha_{e+\frac{1}{2}}$ will be based on a local smoothness indicator which will bias the flux towards the lower order flux $\bm{f}_{e+\frac{1}{2}}$ near regions of lower solution smoothness. However, to enforce admissibility in means (Definition 2), the flux has to be further corrected, as explained in Section 6.3.

### v2

@@ 5.2 Stage 2
$\mathsf{F}^{*}_{e}=a\mathsf{U}^{*}_{e},\qquad\mathsf{U}^{*}_{e}=\mathsf{u}_{e}^{n}-\frac{1}{6}\sigma\mathsf{D}\mathsf{u}_{e}^{n}-\frac{1}{3}\sigma\mathsf{D}\mathsf{u}_{e}^{\ast}=\mathsf{T}^{(2)}\mathsf{u}_{e}^{n}+\mathsf{T}^{(2,\ast)}\mathsf{u}_{e}^{\ast}$
@@
$\mathsf{T}^{(2)}=\mathsf{I}-\frac{1}{6}\sigma\mathsf{D},\qquad\mathsf{T}^{(2,\ast)}=-\frac{1}{3}\sigma\mathsf{D}$
@@
$\mathsf{F}^{*}_{{e+\frac{1}{2}}}=\frac{1}{2}\left[\mathsf{V}_{R}^{T}\mathsf{F}_{e}^{*}+\mathsf{V}_{L}^{T}\mathsf{F}_{e+1}^{*}\right]-\frac{1}{2}a\left(\mathsf{V}_{L}^{T}\mathsf{U}^{*}_{e+1}-\mathsf{V}_{R}^{T}\mathsf{U}^{*}_{e}\right)=a\mathsf{V}_{R}^{T}\mathsf{U}_{e}^{*}\qquad\textrm{and}\qquad\mathsf{F}^{*}_{{e-\frac{1}{2}}}=a\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{*}$
@@
$F_{h}^{\delta}(x_{{e+\frac{1}{2}}}^{-})=\mathsf{V}_{R}^{T}\mathsf{F}_{e}^{*}=a\mathsf{V}_{R}^{T}\mathsf{U}_{e}^{*},\qquad F_{h}^{\delta}(x_{{e-\frac{1}{2}}}^{+})=\text{{{V}}}_{L}^{T}\mathsf{F}_{e}^{*}=a\mathsf{V}_{L}^{T}\mathsf{U}_{e}^{*}$
@@
$\begin{split}F_{{e+\frac{1}{2}}}-F_{h}^{\delta}(x_{{e+\frac{1}{2}}}^{-})&=0\\ F_{{e-\frac{1}{2}}}-F_{h}^{\delta}(x_{{e-\frac{1}{2}}}^{+})&=a\left(\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{*}-\mathsf{V}_{L}^{T}\mathsf{U}_{e}^{*}\right)\end{split}$ (21)
@@
$\begin{split}\partial_{\xi}\mathsf{F}_{h}^{*}&=a\mathsf{D}\mathsf{U}^{*}_{e}+a\mathsf{b}_{L}\left(\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{*}-\mathsf{V}_{L}^{T}\mathsf{U}_{e}^{*}\right)\\ &=a\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{U}_{e-1}^{*}+a\left(\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{T}\right)\mathsf{U}^{*}_{e}\end{split}$ (22)
We now expand $\mathsf{U}_{e}^{*}$ in terms of $\mathsf{u}_{e}^{n}$ as follows
$\mathsf{U}_{e}^{*}=\mathsf{T}^{(2)}\mathsf{u}_{e}^{n}+\mathsf{T}^{(2,\ast)}\mathsf{u}_{e}^{\ast}\qquad\textrm{where}\qquad\mathsf{T}^{(2)}=\mathsf{I}-\frac{1}{6}\sigma\mathsf{D},\qquad\mathsf{T}^{(2,\ast)}=-\frac{1}{3}\sigma\mathsf{D}$
Thus, using the 1-D update formula
$\mathsf{u}^{n+1}=\mathsf{u}^{n}-\frac{\Delta t}{\Delta x_{e}}\partial_{\xi}\mathsf{F}_{h}^{*}$
@@
$\begin{split}\mathsf{A}_{-2}&=-\mathsf{b}_{L}\mathsf{V}_{R}^{T}\mathsf{T}^{(2,\ast)}\mathsf{A}^{(1)}_{-1}\\ \mathsf{A}_{-1}&=\mathsf{b}_{L}\mathsf{V}_{R}^{T}(\mathsf{T}^{(2)}+\mathsf{T}^{(2,\ast)}(1-\sigma\mathsf{A}^{(1)}_{0}))-\sigma(\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{T})\mathsf{T}^{(2,\ast)}\mathsf{A}^{(1)}_{-1}\\ \mathsf{A}_{0}&=-(\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{T})(\mathsf{T}^{(2)}+\mathsf{T}^{(2,\ast)}(\mathsf{I}-\sigma\mathsf{A}^{(1)}_{0}))\\ \mathsf{A}_{+1}&=\mathsf{A}_{+2}=0\end{split}$ (23)
where $\mathsf{A}^{(1)}_{i}$ are defined in (20). The upwind character of D2 flux is reason why we have $\mathsf{A}_{+1}=\mathsf{A}_{+2}=0$.
@@
$\begin{gathered}\widehat{\mathsf{u}}_{k}^{n+1}=H(\sigma,k)\widehat{\mathsf{u}}_{k}^{n}\\ \text{{{H}}}=-\sigma^{2}\mathsf{A}_{-2}\exp(-2i\kappa)-\sigma\mathsf{A}_{-1}\exp(-i\kappa)+\mathsf{I}-\sigma\mathsf{A}_{0}-\sigma\mathsf{A}_{+1}\exp(i\kappa)-\sigma^{2}\mathsf{A}_{+2}\exp(2i\kappa)\end{gathered}$ (24)
@@
The values of $\sigma$ are also sampled in some interval $[\sigma_{\min},\sigma_{\max}]$ and the largest value of $\sigma_{l}\in[\sigma_{\min},\sigma_{\max}]$ for which $\lambda(\sigma_{l})\leq 1$ is determined in a Python code. We start with a large interval $[\sigma_{\min},\sigma_{\max}]$ and then progressively reduce the size of this interval so that the CFL number is determined to about three decimal places. The results of this numerical investigation of stability are shown in Table 1 for two correction functions with polynomial degree $N=3$. The comparison is made with CFL numbers of MDRK-D1 (14) which are experimentally obtained from the linear advection test case (Section 7.1.1), i.e., using a time step size that is slightly larger than these numbers causes the solution to blow up.
@@
LW-D2
$(N=3)$
$\frac{\text{MDRK-D2}}{\text{LW-D2}}$
Radau $\approx$0.09 0.107 1.19 0.103 1.04
$g_{2}$approx0.16 0.224 1.4 0.170 1.31
@@
We see that dissipation model D2 has a higher CFL number compared to dissipation model D1. The CFL numbers for the $g_{2}$ correction function are also significantly higher than those for the Radau correction function. The MDRK scheme also has higher CFL numbers than the single stage LW method [3] for degree $N=3$, which especially with the $g_{2}$ correction function. The optimality of these CFL numbers has been verified by experiments on the linear advection test case (Section 7.1.1), i.e., the solution eventually blows up if the time step is slightly higher than what is allowed by the CFL condition.
@@ 6 Blending scheme
The MDRK scheme (8) gives a high (fourth) order method for smooth problems, but most practical problems involving hyperbolic conservation laws consist of non-smooth solutions like shocks. In such situations, using a higher order method is bound to produce Gibbs oscillations, as stated in Godunov’s order barrier theorem [14]. The cure is to non-linearly add dissipation in regions where the solution is non-smooth, with methods like artificial viscosity, limiters and switching to a robust lower order scheme; the resultant scheme will be non-linear even for linear equations. In this work, we develop a subcell based blending scheme for MDRK similar to the one in [4] for the high order single stage Lax-Wendroff methods. The limiter is applied to each MDRK stage.
@@
$\mathsf{u}^{H,\ast}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t/2}{\Delta x_{e}}\mathsf{R}^{H}_{e},\qquad\mathsf{u}^{H,n+1}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t}{\Delta x_{e}}\mathsf{R}^{\ast,H}_{e}$ (25)
@@
$\mathsf{u}^{L,\ast}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t/2}{\Delta x_{e}}\mathsf{R}^{L}_{e},\qquad\mathsf{u}^{L,n+1}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t}{\Delta x_{e}}\mathsf{R}^{\ast,L}_{e}$ (26)
@@
$\begin{split}\mathsf{u}^{\ast}_{e}&=(1-\alpha_{e})\mathsf{u}^{H,\ast}_{e}+\alpha_{e}\mathsf{u}^{L,\ast}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t/2}{\Delta x_{e}}[(1-\alpha_{e})\mathsf{R}^{H}_{e}+\alpha_{e}\mathsf{R}^{L}_{e}]\\ \mathsf{u}^{n+1}_{e}&=(1-\alpha_{e})\mathsf{u}^{H,n+1}_{e}+\alpha_{e}\mathsf{u}^{L,n+1}_{e}=\mathsf{u}^{n}_{e}-\frac{\Delta t}{\Delta x_{e}}[(1-\alpha_{e})\mathsf{R}^{\ast,H}_{e}+\alpha_{e}\mathsf{R}^{\ast,L}_{e}]\end{split}$ (27)
where $\alpha_{e}\in[0,1]$ must be chosen based on the local smoothness indicator. If $\alpha_{e}=0$ then we obtain the high order MDRK scheme, while if $\alpha_{e}=1$ then the scheme becomes the low order scheme that is less oscillatory. In subsequent sections, we explain the details of the lower order scheme and the design of smoothness indicators. The lower order scheme will either be a first order finite volume scheme or a high resolution scheme based on MUSCL-Hancock idea. In either case, the common structure of the low order scheme can be explained as follows. In the first stage, the lower order residual $\mathsf{R}^{\ast,L}_{e}$ performs evolution from time $t^{n}$ to $t^{n+\frac{1}{2}}$ while, in the second stage, $\mathsf{R}^{L}_{e}$ performs evolution from $t^{n}$ to $t^{n+1}$. It may seem more natural to evolve from $t^{n+\frac{1}{2}}$ to $t^{n+1}$ in the next stage, but that will violate the conservation property because of the expression of second stage of MDRK (625).
@@
$\begin{split}\bm{u}_{e,0}^{L,n+1}&=\bm{u}_{e,0}^{n}-\frac{\Delta t}{w_{0}\Delta x_{e}}[\bm{f}_{\frac{1}{2}}^{e}-\bm{F}^{\ast}_{{e-\frac{1}{2}}}]\\ \bm{u}_{e,p}^{L,n+1}&=\bm{u}_{e,p}^{n}-\frac{\Delta t}{w_{p}\Delta x_{e}}[\bm{f}_{p+\frac{1}{2}}^{e}-\bm{f}_{p-\frac{1}{2}}^{e}],\qquad 1\leq p\leq N-1\\ \bm{u}_{e,N}^{L,n+1}&=\bm{u}_{e,N}^{n}-\frac{\Delta t}{w_{N}\Delta x_{e}}[\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{f}_{N-\frac{1}{2}}^{e}]\end{split}$ (29)
The inter-element fluxes $\bm{F}^{\ast}_{{e+\frac{1}{2}}}$ used in the low order scheme are same as those used in the high order MDRK scheme in equation (7). Usually, Rusanov’s flux [25] will be used for the inter-element fluxes and in the lower order scheme. The element mean value obtained by the low order scheme satisfies
$\overline{\bm{u}}_{e}^{L,n+1}=\sum_{p=0}^{N}\bm{u}_{e,p}^{L,n+1}w_{p}=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast}_{{e-\frac{1}{2}}})$ (30)
@@
$=(1-\alpha_{e})\left[\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast}_{{e-\frac{1}{2}}})\right]+\alpha_{e}\left[\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast}_{{e-\frac{1}{2}}})\right]$
$=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast}_{{e-\frac{1}{2}}})$
@@
$\begin{split}\overline{\bm{u}}^{\ast}_{e}&=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t/2}{\Delta x_{e}}(\bm{F}_{{e+\frac{1}{2}}}-\bm{F}_{{e-\frac{1}{2}}})\\ \overline{\bm{u}}_{e}^{n+1}&=\overline{\bm{u}}_{e}^{n}-\frac{\Delta t}{\Delta x_{e}}(\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast}_{{e-\frac{1}{2}}})\end{split}$ (31)
@@
The inter-element fluxes $\bm{F}_{{e+\frac{1}{2}}},\bm{F}^{\ast}_{{e+\frac{1}{2}}}$ are used both in the low and high order schemes. To achieve high order accuracy in smooth regions, this flux needs to be high order accurate, however it may produce spurious oscillations near discontinuities when used in the low order scheme. A natural choice to balance accuracy and oscillations is to take
$\begin{split}\bm{F}_{{e+\frac{1}{2}}}&=(1-\alpha_{{e+\frac{1}{2}}})\bm{F}_{{e+\frac{1}{2}}}^{\text{HO}}+\alpha_{{e+\frac{1}{2}}}\bm{f}_{{e+\frac{1}{2}}}\\ \bm{F}^{\ast}_{{e+\frac{1}{2}}}&=(1-\alpha_{{e+\frac{1}{2}}})\bm{F}^{\text{HO}\ast}_{{e+\frac{1}{2}}}+\alpha_{{e+\frac{1}{2}}}\bm{f}_{{e+\frac{1}{2}}}\end{split}$ (32)
where $\bm{F}^{\text{HO}}_{{e+\frac{1}{2}}},\bm{F}^{\text{HO}\ast}_{{e+\frac{1}{2}}}$ are the high order inter-element time-averaged numerical fluxes used in the MDRK scheme (15) and $\bm{f}_{{e+\frac{1}{2}}}$ is a lower order flux at the face $x_{{e+\frac{1}{2}}}$ shared between FR elements and subcells (3334). The blending coefficient $\alpha_{{e+\frac{1}{2}}}$ will be based on a local smoothness indicator which will bias the flux towards the lower order flux $\bm{f}_{{e+\frac{1}{2}}}$ near regions of lower solution smoothness. However, to enforce admissibility in means (Definition 2), the flux has to be further corrected, as explained in Section 6.3.